OpenStudy (anonymous):

MEDAL AND FAN :D Jacob deposits $60 into an investment account with an interest rate of 4%, compounded annually. The equation, y = 60(1 + 0.04)x, can be used to determine the number of years, x, it takes for Jacob's balance to reach a certain amount of money, y. Jacob graphs the relationship between time and money. 1. What is the y-intercept of Jacob's graph? 2. What is the equation of the asymptote? 3. If Jacob doesn't deposit any additional money into the account, how much money will he have in eight years? Round your answer to the nearest cent.

OpenStudy (shamim):

i think ur equation will b y=60(1+0.04)^x

OpenStudy (shamim):

i wanna try in my own way. i think we hv to take ln of both sides

OpenStudy (shamim):

ln y=ln (60*(1.04)^x)

OpenStudy (shamim):

r u following me?

OpenStudy (shamim):

plz feel free to ask any question

OpenStudy (shamim):

ln y=ln 60+ln (1.04)^x

OpenStudy (shamim):

ln y=x*ln 1.04+ln 60

OpenStudy (shamim):

frm above equation can u tell me the y intetcep